OpenStudy (anonymous):

differentiate f(u)= ln(u^2-1)^3 =(1/((u^2-1)^3)))(6u(u^2-1)^2) =(6u(u^2-1)^2)/(u^2-1)^3 =6u/(u^2-1)^3 is that correct?

OpenStudy (campbell_st):

nope... you seem to have made an error on cancelling you have in the 2nd line of working \[=\frac{6u(u^2 -1)^2}{(u^2-1)^3}\] if you remove a common factor of \[(u^2 -1)^2 \] you would be left with \[=\frac{6u}{u^2-1}\] hope this helps

OpenStudy (anonymous):

thank you so much